# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

### v2

@@ 1 Introduction
Music recordings produced in amateur settings often suffer from a variety of quality issues that distinguish them from professionally mastered recordings (Wilson & Fazenda, 2016; Mourgela et al., 2024; Deruty & Tardieu, 2014). For instance, an enthusiast recording vocals in a garage may introduce excessive reverberation, making the voice sound distant and “echoey.” Similarly, using inexpensive microphones or misconfigured interfaces can lead to distortion and clipping when loud peaks exceed the recording range, resulting in harsh crackles or flattened dynamics (Zang et al., 2025). Tonal imbalances are also common: a home recording might sound overly “muddy” or “tinny” if certain frequency bands dominate or vanish due to poor room acoustics or improper microphone placement. Even the stereo image can be narrowed or skewed, reducing the sense of space in the mix. In practice, engineers address these problems with specialized tools: e.g., dereverberation plugins to remove room echo, declipping algorithms to reconstruct saturated peaks, equalizers to rebalance frequencies, and stereo enhancers to widen the image. Mastering a flawed track has become a labor-intensive process requiring expert skill and multiple stages of manual adjustment.
@@
The need for an automated all-in-one solution is evident. Creators with limited resources often lack the expertise to apply the right combination of restoration tools, and a piecemeal approach may fail to fully recover a track’s fidelity. This motivates SonicMaster, a unified approach to music restoration and mastering that can correct a broad spectrum of audio degradations within a single model. We introduce a single flow-based generative framework (Liu et al., 2022; Esser et al., 2024) that simultaneously performs dereverberation, equalization, declipping, dynamic-range expansion, and stereo enhancement. The backbone is trained on a curated corpus of polyphonic music rendered through a combinatorial grid of simulated degradations, enabling the network to learn the joint statistics and cross-couplings of common artifacts rather than treating them in isolation. This joint training eliminates the need for error-prone cascades of task-specific modules and reduces inference to a single forward pass.
@@
Crucially, SonicMaster incorporates multimodal conditioning through natural language instructions that capture production objectives. A prompt such as reduce the hollow room sound attenuates late reflections without suppressing desirable early reverberation, whereas increase the brightness selectively enhances the treble frequencies while preserving spectral balance elsewhere.In the absence of a prompt, SonicMaster switches to an automatic mode that applies perceptually balanced mastering. Existing speech restoration models (e.g. VoiceFixer (Liu et al., 2021)) also address artifacts sequentially, ignoring their mutual influence. By unifying restoration and mastering tasks under a single, prompt-driven generative model, SonicMaster delivers professional-grade improvements while affording fine-grained creative control. Recent advances such as Mustango (Text-guided music generation) (Melechovsky et al., 2024), FlowSep (Yuan et al., 2025) (text-guided source separation), TangoFlux (Hung et al., 2024) (reward-optimized text-to-audio diffusion), and instruction-guided models like AUDIT (Wang et al., 2023) or AudioLDM/AudioLDM2 (Liu et al., 2023; 2024) illustrate powerful generative methods, but they target orthogonal tasks—generation, separation, or localized editing—rather than unified restoration. In contrast, SonicMaster uniquely addresses comprehensive multi-artifact music restoration and mastering through a single controllable rectified-flow architecture, bridging dereverberation, declipping, tonal rebalancing, dynamics, and stereo enhancement under prompt guidance.
@@ 2 Related Work
Restoring and mastering audio spans speech and music enhancement, audio inpainting, and source separation—areas that have mostly been handled separately (Záviška et al., 2020). Diffusion-based generative models and text-guided audio editing (Hou et al., 2025; Jiang et al., 2025; Zhang et al., 2024; Manor & Michaeli, 2024) now let us tackle these problems together. We review these advances, their uses, and the gaps that SonicMaster aims to fill. Early audio restoration efforts typically focused on single domains or isolated tasks, addressing issues like noise, clipping, or reverb in separation. Speech enhancement (Yousif & Mahmmod, 2025) and music enhancement evolved largely independently, and tasks such as audio inpainting or source separation were treated with specialized methods (Lemercier et al., 2025).
@@
Early signal-model and interpolation methods could patch only very short gaps (<10<10 ms), leaving longer dropouts unresolved. Deep generative models now bridge that gap: diffusion-based systems convincingly regenerate missing music sections and clipped peaks (Moliner & Välimäki, 2023). The authors in (Wang et al., 2023; Liu et al., 2023) extend this with instruction-guided diffusion for audio inpainting, while VoiceFixer (Liu et al., 2021) jointly denoises, dereverbs, and declipse speech, though it is restricted to voice and does not offer user control. In music, (Imort et al., 2022) removed heavy guitar distortion (including clipping) with neural networks, surpassing sparse-optimization baselines in quality and speed. (Lee et al., 2024) introduce a pruning approach to recover sparse audio effect chains from mixed recordings, essentially reverse-engineering mixing graphs from input/output pairs. Alongside works like (Bhandari et al., 2025) on iterative corruption refinement, (Steinmetz et al., 2021) on differentiable mixing consoles and (Martínez-Ramírez et al., 2022) on out-of-domain mixing generalization, and diffusion restorers such as MaskSR (Li et al., 2024), these highlight emerging methods that bridge audio restoration with controllable, interpretable effect modeling.
@@
Research on learning-based equalization is still emerging. In recently proposed work (Mockenhaupt et al., 2024), the authors introduce CNN-based approach to automatically equalize instrument stems by predicting parametric EQ settings, showing improvements over earlier heuristic methods. Notably, the VoiceFixer (Liu et al., 2021) addressed bandwidth extension, essentially restoring high-frequency content as part of its speech restoration, which can be seen as a form of equalization correction. Similarly, diffusion-based restorers like MaskSR (Li et al., 2024) treat low-frequency muffling as a distortion to fix, using discrete token prediction to restore a balanced spectrum. In Text2FX (Chu et al., 2025), CLAP (Elizalde et al., 2023) is used in inference mode to steer the parameters of EQ and reverb audio effects.
@@ 3.1 Dataset
In the absence of text-conditioned music restoration datasets, we generate SonicMaster dataset by pairing high-quality audio with systematically applied degradations and corresponding natural language instructions. Our source comprises songs from (Roy et al., 2025) and additional content from Jamendo [^2] https://www.jamendo.com/ under Creative Commons licence using the official Jamendo API. In total, we have sourced 580580k recordings. We ensure balanced genre representation by defining $10$ groups, where each group consists of multiple semantically related genre tags, e.g., Hip-Hop genre group containing the following tags: ”rap”, ”hiphop”, ”trap”, ”alternativehiphop”, ”gangstarap”. Complete taxonomies are provided in the Appendix. Track selection employs Audiobox Aesthetics toolbox (Tjandra et al., 2025) for automated production quality assessment. We select $2,500$ songs per genre groups using adaptive production quality thresholds ranging from $6.5$ to $8$ to balance comprehensive sub-genre representation with sufficient production quality. We extract random 30s excerpts from each track, positioned between 1515% and 8585% of its total duration. The complete pipeline is illustrated in Figure 1. We train SonicMaster by applying $19$ distinct degradations to the audio and pairing each with a matching natural language editing instruction. The degradations span five classes: (i) EQ, (ii) Dynamics, (iii) Reverb, (iv) Amplitude, and (v) Stereo.
Figure 1: SonicMaster dataset creation pipeline and overview.
Equalization (EQ): Spectral degradations cover $10$ effects targeting perceptual audio characteristics: Brightness, Darkness, Airiness, Boominess, Muddiness, Warmth, Vocals, Clarity, Microphone, and X-band. Brightness, Darkness, Airiness, Boominess, and Warmth are emulated with low- or high-shelf EQ; Clarity with a Butterworth low-pass filter; Vocals and Muddiness with Chebyshev-II band-pass filters. Microphone applies one of $20$ Poliphone transfer functions (Salvi et al., 2025), while X-band uses an 881212-band, logarithmically spaced peaking EQ with $\textpm 6$ dB gain per band.
@@
Reverb: The Reverb category contains four distinct approaches: three of them utilise the Pyroomacoustics library (Scheibler et al., 2018), which simulates acoustic environments with the image source method. We simulate three types of rooms: Small, Big, and Mixed. For our fourth Reverb function, we utilize 12 selected room impulse responses from the openAIR library dataset (Howard & Angus, n.d.), which give us audio with more real-life properties. The resulting impulse responses from all the functions are convoluted with the clean signals.
@@
Each ground truth yields $7$ corrupted variants: $4$ with a single, $2$ with double, and $1$ with triple degradation. In multi-degradation, we sample at most one effect from each of the $5$ categories, so an EQ choice, for instance, blocks further EQ picks. To avoid duplicates in the single-degradation set, high-probability effects with narrow parameter ranges: Stereo, Clipping, and Punchare, are used only once. Each degradation is linked to a one-sentence instruction randomly selected from 8–10 expert-written options; these sentences are concatenated into the full prompt, and we store two prompt variants per clip for robustness. We also record every applied effect and its parameters (gain, absorption), supporting tasks such as parameter prediction. For Compression and Reverb, there is a 1515% chance of injecting “hidden clipping” with no corresponding instruction. When neither hidden clipping nor an Amplitude effect is present, the audio is peak-normalised to a random level between 0.81.00.8-1.0. Further details can be found in Appendix.
@@ 3.2 SonicMaster Architecture
SonicMaster employs a hybrid architecture combining Multimodal Diffusion Transformer (MM-DiT)(Esser et al., 2024) blocks with subsequent Diffusion Transformers (DiT) layers (Peebles & Xie, 2023). As outlined in Figure 2, stereo waveforms (44.1 kHz) ($x_{t}$) undergo VAE ecoding (Evans et al., 2024) into compact spectro-temporal latent representations. Restoration, therefore, occurs entirely in this learned space, allowing large receptive fields without sample-level overhead. The MM-DiT processes degraded latent representations alongside the text embeddings from a frozen FLAN-T5 encoder (Chung et al., 2024). The resulting conditioned representations pass through subsequent DiT layers to predict flow velocity $v_{t}$, steering the latent toward its clean target $\hat{x}_{t}$. Prompts like ”reduce reverb” biases this prediction trajectory to suppress decay tails, while the downstream DiT layersrefine musical coherence. A pooled-audio branch, active in 2525% of training cases, concatenates a temporally averaged 5155–15s clean cue with the pooled prompt embedding and injects it at every MM-DiT/DiT layer, enabling seamless chaining of $30$s segments for long-form generation while degrading gracefully when no reference is supplied.
@@
We adopt the Stable Audio Open VAE (Evans et al., 2024) to encode–decode stereo signals sampled at 44.1 kHz, yielding a compact latent representation while retaining high-fidelity reconstruction. Text instructions are embedded with FLAN-T5 Large (Chung et al., 2024); the resulting tensor $c_{\text{text}}\in\mathbb{R}^{B\times S_{\text{text}}\times D_{\text{text}}}$ (with $D_{\text{text}}=1024$) is used as a conditioning signal.
@@
SonicMaster employs rectified flow (Liu et al., 2022; Esser et al., 2024), to predict flow velocity from degraded to clean audio in latent space, unlike other models that map noise to output distributions (Fei et al., 2024; Hung et al., 2024).
@@ 4.1 Baselines and training setup
We train SonicMaster using 5 NVIDIA L40S GPUs for 40 epochs with a total batch size of 80. We adopt classifier-free guidance (Ho & Salimans, 2022) by (i) dropping the text prompt in 1010% of samples and (ii) replacing it in another 1010% with one of four generic phrases (“Make it sound better!”, “Master this track for me, please!”, “Improve this!”, “Can you improve the sound of this song?”). In 2525% of cases, the model is additionally conditioned—via the pooling branch—on the first 10 s of clean audio. Unless stated otherwise, all experiments follow these conditioning settings while comparing multiple SonicMaster variants and baselines.
@@
[…]
@@
Mel2Mel + Diffwave Kandpal et al. (2022) 0.0278 0.3561 0.0049 0.0135 0.0855 0.4705 0.4436 0.256 0.2604 0.1885
Text2FX-EQ (Chu et al., 2025) 0.0219 0.3809 0.0055 0.0276 0.2112 0.3651 0.4955 0.2199 0.4441 0.3419
   -w/o Text Cond. 0.0130 0.1432 0.0032 0.0101 0.0086 0.0448 0.0841 0.0668 0.0154 0.0424
[…]
@@
SonicMasterLarge (6/18) 0.0102 0.0639 0.0021 0.0060 0.0065 0.0329 0.0510 0.0517 0.0070 0.0289
Evaluation is conducted along two orthogonal axes. (i) Global perceptual fidelity is quantified with FAD on CLAP embeddings (Elizalde et al., 2023), Kullback–Leibler divergence (KL), structural similarity (SSIM) on 128-bin mel-spectrograms, and the Production Quality (PQ) score from the Audiobox Aesthetics toolbox (Tjandra et al., 2025). (ii) Degradation-specific restoration efficacy is measured by average absolute error reduction: for every degraded clip in a $7000$ clip test set, we compute the relevant (based on the degradation deployed) artefact-aware metric against its clean counterpart from a $1000$ sample reference set, then recompute the metric after SonicMaster processing; the relative decrease indicates how closely each model variant approaches the ground-truth.
For X-band EQ and microphone-TF degradations, we compute the spectral balance over eight bands and report their cosine distance. All other EQ effects are scored by the energy ratio between the affected band and the full spectrum. Compression is measured as the standard deviation of frame-level RMS (2048-sample frames, 1024 hop); punch as the mean onset-envelope value (librosa.onset.onset_strength). Because RT60 estimates are unreliable on dense mixes, reverb is assessed via the Euclidean distance of modulation spectra. Clipping uses spectral flatness; volume, the global RMS; and stereo width, the RMS ratio of the mid and side signals, $\mathrm{RMS}\!\left[\frac{L-R}{2}\right]/\mathrm{RMS}\!\left[\frac{L+R}{2}\right]$.
We presented listeners with $43$ audio sample pairs – degraded inputs and SonicMaster outputs – to rate, consisting of $2$ pairs for each degradation function ($2\times 19=38$ single degraded samples), $3$ pairs of double and $2$ pairs of triple degraded samples. Using a $7$-point Likert Scale, listeners were to rate: 1) The extent of improvement from the input to SonicMaster output represented by the text prompt (Text relevance), 2) audio quality of input (Quality1), 3) audio quality of the inferred SonicMaster sample (Quality2), 4) consistency and fluency of the inferred sample (Consistency), and 5) preference between the two samples, where 1 represents full preference of the ground truth degraded input, and 7 represents the SonicMaster inferred sample (Preference). The study was attended by 8 listeners (5 music experts and 3 Music Information Retrieval researchers).
Furthermore, to benchmark against existing methods, we conducted an additional study with 13 participants comparing SonicMaster against Text2FX (Chu et al., 2025) and Mel2Mel + Diffwave (Kandpal et al., 2022) on 20 randomly selected samples from our test set. The evaluation included 10 samples with X-band EQ degradation and 10 with reverberation artifacts. Note that Text2FX is limited to EQ effects (as its reverb effect is only additive, thus excluded), making direct comparison feasible only for the EQ category. Since the baseline methods’ evaluation sets are not publicly available, we performed this comparison exclusively on our curated test data.
5 Results
5.1 Objective Evaluation
Degradation-Specific Performance: Tables 4.1 and 5.1 demonstrates SonicMaster’s superiority over baselines of Text2FX in EQ, and WPE/HPSS in Reverb. SonicMaster improves in all categories when compared to the degraded and reconstructed inputs. Furthermore, the reconstructed input metrics are overall slightly worse (with exceptions) than those of the ground truth degraded inputs.
Model (MMDiT/DiT) Reverb Dynamics Amplitude Stereo
Small Big Mix Real Compressor Punch Clip Volume
[…]
@@ 5.1 Objective Evaluation
WPE (Nakatani et al., 2010) 0.4849 0.4732 0.5207 0.4854 - - - - -
Mel2Mel + Diffwave (Kandpal et al., 2022) 0.4404 0.4387 0.4361 0.4368 - - - - -
   -w/o Text Cond. 0.3732 0.3805 0.4012 0.3264 0.0157 0.0730 2.812 0.0465 0.1416
@@
SonicMasterLarge (6/18) 0.3954 0.4511 0.4191 0.4066 0.0258 0.1101 3.734 0.0424 0.0850
Table 2: Objective scores: Reverb, Dynamics, Amplitude, and Stereo. Clip are multiplied by 1000.
Model Single deg. Double+triple deg. All
[…]
Ground truth mastered reference - - - 7.886 - - - 7.886 - - - 7.886
@@
Mel2Mel + Diffwave (Kandpal et al., 2022) 0.522 14.938 0.4465 6.158 0.474 15.185 0.416 5.953 0.491 15.044 0.433 6.070
   -w/o Text Cond. 0.069 0.917 0.621 7.772 0.088 1.484 0.586 7.643 0.074 1.160 0.606 7.716
[…]
Ground truth mastered reference - - - 7.885 - - - 7.885 - - - 7.885

### v3

@@ 1 Introduction
Music recordings produced in amateur settings often suffer from a variety of quality issues that distinguish them from professionally mastered recordings (wilson2016perception; mourgela2024exploring; deruty2014dynamic). For instance, an enthusiast recording vocals in a garage may introduce excessive reverberation, making the voice sound distant and “echoey.” Similarly, using inexpensive microphones or misconfigured interfaces can lead to distortion and clipping when loud peaks exceed the recording range, resulting in harsh crackles or flattened dynamics (zang2025music). Tonal imbalances are also common: a home recording might sound overly “muddy” or “tinny” if certain frequency bands dominate or vanish due to poor room acoustics or improper microphone placement. Even the stereo image can be narrowed or skewed, reducing the sense of space in the mix. In practice, engineers address these problems with specialized tools: e.g., dereverberation plugins to remove room echo, declipping algorithms to reconstruct saturated peaks, equalizers to rebalance frequencies, and stereo enhancers to widen the image. Mastering a flawed track has become a labor-intensive process requiring expert skill and multiple stages of manual adjustment.
@@
The need for an automated all-in-one solution is evident. Creators with limited resources often lack the expertise to apply the right combination of restoration tools, and a piecemeal approach may fail to fully recover a track’s fidelity. This motivates SonicMaster, a unified approach to music restoration and mastering that can correct a broad spectrum of audio degradations within a single model. We introduce a single flow-based generative framework (liu2022flow; esser2024scaling) that simultaneously performs dereverberation, equalization, declipping, dynamic-range expansion, and stereo enhancement. The backbone is trained on a curated corpus of polyphonic music rendered through a combinatorial grid of simulated degradations, enabling the network to learn the joint statistics and cross-couplings of common artifacts rather than treating them in isolation. This joint training eliminates the need for error-prone cascades of task-specific modules and reduces inference to a single forward pass.
@@
Crucially, SonicMaster incorporates multimodal conditioning through natural language instructions that capture production objectives. A prompt such as reduce the hollow room sound attenuates late reflections without suppressing desirable early reverberation, whereas increase the brightness selectively enhances the treble frequencies while preserving spectral balance elsewhere.In the absence of a prompt, SonicMaster switches to an automatic mode that applies perceptually balanced mastering. Existing speech restoration models (e.g. VoiceFixer by liu2021voicefixer) also address artifacts sequentially, ignoring their mutual influence. By unifying restoration and mastering tasks under a single, prompt-driven generative model, SonicMaster delivers professional-grade improvements while affording fine-grained creative control. Recent advances such as Mustango (Text-guided music generation) by melechovsky2024mustango, FlowSep by yuan2025flowsep (text-guided source separation), TangoFlux by hung2024tangoflux (reward-optimized text-to-audio diffusion), and instruction-guided models like AUDIT (wang2023audit) or AudioLDM/AudioLDM2 (liu2023audioldm; liu2024audioldm) illustrate powerful generative methods, but they target orthogonal tasks—generation, separation, or localized editing—rather than unified restoration. In contrast, SonicMaster uniquely addresses comprehensive multi-artifact music restoration and mastering through a single controllable rectified-flow architecture, bridging dereverberation, declipping, tonal rebalancing, dynamics, and stereo enhancement under prompt guidance.
@@
Figure 1: SonicMaster dataset creation pipeline and overview.
@@ 2 Related Work
Restoring and mastering audio spans speech and music enhancement, audio inpainting, and source separation—areas that have mostly been handled separately zavivska2020survey. Diffusion-based generative models and text-guided audio editing 10890309; jiang2025listen; zhang2024musicmagus; manor2024zero; haninstructme now let us tackle these problems together. We review these advances, their uses, and the gaps that SonicMaster aims to fill. Early audio restoration efforts typically focused on single domains or isolated tasks, addressing issues like noise, clipping, or reverb in separation. Speech enhancement yousif2025speech and music enhancement evolved largely independently, and tasks such as audio inpainting or source separation were treated with specialized methods lemercier2025diffusion.
@@
Early signal-model and interpolation methods could patch only very short gaps (<10<10 ms), leaving longer dropouts unresolved. Deep generative models now bridge that gap: diffusion-based systems convincingly regenerate missing music sections and clipped peaks moliner2023diffusion. The authors in wang2023audit; liu2023audioldm extend this with instruction-guided diffusion for audio inpainting, while VoiceFixer liu2021voicefixer jointly denoises, dereverbs, and declipse speech, though it is restricted to voice and does not offer user control. In music, imort2022distortion removed heavy guitar distortion (including clipping) with neural networks, surpassing sparse-optimization baselines in quality and speed. lee2024searching introduce a pruning approach to recover sparse audio effect chains from mixed recordings, essentially reverse-engineering mixing graphs from input/output pairs. Alongside works like bhandari2025improvnet on iterative corruption refinement, steinmetz2021automatic on differentiable mixing consoles and martinez2022automatic on out-of-domain mixing generalization, and diffusion restorers such as MaskSR li2024masksr, these highlight emerging methods that bridge audio restoration with controllable, interpretable effect modeling. Moreover, rice2023general introduce a compositional architecture for multi-effect audio removal using effect-specific removal modules.
@@
Research on learning-based equalization is still emerging. mockenhaupt2024automatic recently introduced CNN-based approach to automatically equalize instrument stems by predicting parametric EQ settings, showing improvements over earlier heuristic methods. Notably, the VoiceFixer (liu2021voicefixer) addressed bandwidth extension, essentially restoring high-frequency content as part of its speech restoration, which can be seen as a form of equalization correction. Similarly, diffusion-based restorers like MaskSR li2024masksr treat low-frequency muffling as a distortion to fix, using discrete token prediction to restore a balanced spectrum. In Text2FX chu2025text2fx, CLAP elizalde2023clap is used in inference mode to steer the parameters of EQ and reverb audio effects.
@@ 3.1 Dataset
In the absence of text-conditioned music restoration datasets, we generate SonicMaster dataset by pairing high-quality audio with systematically applied degradations and corresponding natural language instructions. Our source comprises songs from roy2025jamendomaxcaps and additional content from Jamendo [^2] https://www.jamendo.com/ under Creative Commons licence using the official Jamendo API. In total, we have sourced 580580k recordings. We ensure balanced genre representation by defining $10$ groups, where each group consists of multiple semantically related genre tags, e.g., Hip-Hop genre group containing the following tags: “rap”, “hiphop”, “trap”, “alternativehiphop”, “gangstarap”. Complete taxonomies are provided in the Appendix. Track selection employs Audiobox Aesthetics toolbox (tjandra2025meta) for automated production quality assessment. We select $2,500$ songs per genre groups using adaptive production quality thresholds ranging from $6.5$ to $8$ to balance comprehensive sub-genre representation with sufficient production quality. We extract random 30s excerpts from each track, positioned between 1515% and 8585% of its total duration. The complete pipeline is illustrated in Figure 1. We train SonicMaster by applying $19$ distinct degradations to the audio and pairing each with a matching natural language editing instruction. The degradations span five classes: (i) EQ, (ii) Dynamics, (iii) Reverb, (iv) Amplitude, and (v) Stereo.
Equalization (EQ): Spectral degradations cover $10$ effects targeting perceptual audio characteristics: Brightness, Darkness, Airiness, Boominess, Muddiness, Warmth, Vocals, Clarity, Microphone, and X-band. Brightness, Darkness, Airiness, Boominess, and Warmth are emulated with low- or high-shelf EQ; Clarity with a Butterworth low-pass filter; Vocals and Muddiness with Chebyshev-II band-pass filters. Microphone applies one of $20$ Poliphone transfer functions (salvi2025poliphone), while X-band uses an 881212-band, logarithmically spaced peaking EQ with $\textpm 6$ dB gain per band.
@@
Reverb: The Reverb category contains four distinct approaches: three of them utilise the Pyroomacoustics library (scheibler2018pyroomacoustics), which simulates acoustic environments with the image source method. We simulate three types of rooms: Small, Big, and Mixed. For our fourth Reverb function, we utilize 12 selected room impulse responses from the openAIR library dataset (openairlibrary), which give us audio with more real-life properties. The resulting impulse responses from all the functions are convoluted with the clean signals.
@@
Each ground truth yields $7$ corrupted variants: $4$ with a single, $2$ with double, and 11 with triple degradation. In multi-degradation, we sample at most one effect from each of the $5$ categories, so an EQ choice, for instance, blocks further EQ picks. To avoid duplicates in the single-degradation set, high-probability effects with narrow parameter ranges: Stereo, Clipping, and Punch, are used only once (in the 4 versions per original, e.g., there cannot be two single-degraded versions with Stereo degradation, as they would be identical). Each degradation is linked to a one-sentence instruction from 8–10 possible options (all written by a music expert); these sentences are concatenated into the full prompt, and we store two prompt variants per clip for robustness. We also record every applied effect and its parameters (gain, absorption), supporting tasks such as parameter prediction. For Compression and Reverb, there is a 1515% chance of injecting “hidden clipping” with no corresponding instruction to emulate real life cases of constructive interference in a reverberant room, or overcompensated gain setting of a compressor. When neither hidden clipping nor an Amplitude effect is present, the audio is peak-normalised to a random level between 0.81.00.8-1.0. Further details can be found in Appendix.
@@ 3.2 SonicMaster Architecture
SonicMaster employs a hybrid architecture combining Multimodal Diffusion Transformer (MM-DiT)(esser2024scaling) blocks with subsequent Diffusion Transformers (DiT) layers (peebles2023scalable). As outlined in Figure 2, stereo waveforms (44.1 kHz) ($x_{t}$) undergo VAE ecoding (evans2024fast) into compact spectro-temporal latent representations. Restoration, therefore, occurs entirely in this learned space, allowing large receptive fields without sample-level overhead. The MM-DiT processes degraded latent representations alongside the text embeddings from a frozen FLAN-T5 encoder (chung2024scaling). The resulting conditioned representations pass through subsequent DiT layers to predict flow velocity $v_{t}$, steering the latent toward its clean target $\hat{x}_{t}$. Prompts like “reduce reverb” biases this prediction trajectory to suppress decay tails, while the downstream DiT layers refine musical coherence. A pooled-audio branch, active in 2525% of training cases, concatenates a temporally averaged 5155–15s clean cue with the pooled prompt embedding and injects it at every MM-DiT/DiT layer, enabling seamless chaining of $30$s segments for long-form generation while degrading gracefully when no reference is supplied.
@@
We adopt the Stable Audio Open VAE (evans2024fast) to encode–decode stereo signals sampled at 44.1 kHz, yielding a compact latent representation while retaining high-fidelity reconstruction. Text instructions are embedded with FLAN-T5 Large (chung2024scaling); the resulting tensor $c_{\text{text}}\in\mathbb{R}^{B\times S_{\text{text}}\times D_{\text{text}}}$ (with $D_{\text{text}}=1024$) is used as a conditioning signal.
@@
SonicMaster employs rectified flow (liu2022flow; esser2024scaling), to predict flow velocity from degraded to clean audio in latent space, unlike other models that map noise to output distributions (fei2024flux; hung2024tangoflux).
@@ 4.1 Baselines and training setup
We train SonicMaster using 5 NVIDIA L40S GPUs for 40 epochs with a total batch size of 80. We adopt classifier-free guidance (ho2022classifier) by (i) dropping the text prompt in 1010% of samples and (ii) replacing it in another 1010% with one of four generic phrases (“Make it sound better!”, “Master this track for me, please!”, “Improve this!”, “Can you improve the sound of this song?”). In 2525% of cases, the model is additionally conditioned—via the pooling branch—on the first 10 s of clean audio. Unless stated otherwise, all experiments follow these conditioning settings while comparing multiple SonicMaster variants and baselines.
@@
Given that Text2FX [^3] Appendix A.3 has details of the Text2FX-directional, both loss formulation and EQ prompt construction. is not a restoration model, we further deploy its directional variant as a meaningful text-guided audio manipulation baseline. SonicMaster operates in a text-conditioned enhancement paradigm, where the model must follow natural-language instructions (e.g., “reduce muddiness”, “increase clarity”). Text2FX-directional is specifically designed for instruction-following tasks: it steers the audio embedding in the same semantic direction defined by a target prompt and its contrast prompt.
Model Clarity Boom Airy Bright Dark Muddy Warm Vocals Mic. X-band
Snippet Evaluation (Short Segments)
@@
Mel2Mel + Diffwave kandpal2022music 0.0278 0.3561 0.0049 0.0135 0.0855 0.4705 0.4436 0.2560 0.2604 0.1885
Text2FXcos chu2025text2fx 0.0219 0.3809 0.0055 0.0276 0.2112 0.3651 0.4955 0.2199 0.4441 0.3419
Text2FXdir chu2025text2fx 0.0421 0.3977 0.0206 0.0143 0.3021 0.2602 0.5461 0.2517 0.6120 0.5038
SonicMaster (Ours) 0.0114 0.0834 0.0019 0.0059 0.0058 0.0388 0.0617 0.0576 0.0088 0.0358
Full Song Evaluation (Long-Form)
Ablation – No Text Condition 0.0130 0.1432 0.0032 0.0101 0.0086 0.0448 0.0841 0.0668 0.0154 0.0424
Ablation – Shuffled Prompts 0.0187 0.2075 0.0077 0.0132 0.0362 0.0981 0.1648 0.1043 0.0424 0.0998
Full Song Evaluation (Long-Form)
@@
SonicMaster (Ours) 0.0102 0.0639 0.0021 0.0060 0.0065 0.0329 0.0510 0.0517 0.0070 0.0289
Table 1: EQ Objective Evaluation (Average Absolute Error). Bold = best performance (lowest error). SonicMaster outperforms baselines in all categories in snippet and full-song scenarios.
Model Reverb Dynamics Amplitude Stereo
Small Big Mix Real Comp. Punch Clip Vol.
Snippet Evaluation (Short Segments)
@@
WPE nakatani2010speech 0.4849 0.4732 0.5207 0.4854 - - - - -
Mel2Mel + Diffwave kandpal2022music 0.4404 0.4387 0.4361 0.4368 - - - - -
SonicMaster (Ours) 0.3663 0.3726 0.3935 0.3109 0.0193 0.0871 1.506 0.0468 0.1058
Ablation Studies
Ablation – No Text Condition 0.3732 0.3805 0.4012 0.3264 0.0157 0.0730 2.812 0.0465 0.1416
Ablation – Shuffled Prompts 0.4161 0.4236 0.4538 0.3903 0.0225 0.0895 2.874 0.0895 0.3213
Full Song Evaluation (Long-Form)
SonicMaster (Ours) 0.3954 0.4511 0.4191 0.4066 0.0258 0.1101 3.734 0.0424 0.0850
Table 2: Objective Scores: Reverb, Dynamics, Amplitude, and Stereo. Clip scores are multiplied by 1000. Bold indicates best performance (lowest error).
Model Single Deg. Double+Triple Deg. All
@@
Snippet Evaluation (Short Segments)
GT Mastered Ref. - - - 7.886 - - - 7.886 - - - 7.886
@@
Mel2Mel + Diffwave kandpal2022music 0.522 14.938 0.447 6.158 0.474 15.185 0.416 5.953 0.491 15.044 0.433 6.070
SonicMaster (Ours) 0.069 0.696 0.624 7.743 0.082 1.145 0.589 7.654 0.073 0.888 0.609 7.705
Ablation Studies
Ablation – No Text Condition 0.069 0.917 0.621 7.772 0.088 1.484 0.586 7.643 0.074 1.160 0.606 7.716
Ablation – Shuffled Prompts 0.081 2.014 0.598 7.610 0.131 3.249 0.558 7.283 0.098 2.543 0.581 7.470
Full Song Evaluation (Long-Form)
GT Mastered Ref. - - - 7.885 - - - 7.885 - - - 7.885